OpenStudy (anonymous):

_____ Evaluate √16-x when x = 8. Write the answer as a decimal. Round to the nearest hundredth if necessary.

sam (.sam.):

\[\sqrt{16-x} , x=8\] \[\sqrt{16-8}\] \[\sqrt{8}\] \[\approx 2.83\]